Prev Next

SW&IT Meeting Minutes 2022-02-14

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220207
LR20220207 - Bec, Matthieu 40.0
LR20220207 - Mechtley, Brandon 0.0
LR20220207 - Eychaner, Glenn 0.0
LR20220207 - Edmeades, David 40.0
LR20220207 - Grenz, Paul 0.0
LR20220207 - Hooper, Stephen 0.0
LR20220207 - Zhang, Yang 40.0
LR20220207 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • OC, Management, meet Joseph Shields
    • ATT PM
    • archive meeting, LBTI data plans
    • set_z0 follow-up
    • DMS archive mongo, lit2/webpack
    • PB work preps readiness, JPH did nice lab setup
    • shark-nir rfa's

  • This week
    • set_z0 follow-up
    • all-hands presentation ???-??
    • DMS archive mongo, lit2/webpack ctd., investigate SSE hooks
    • shark-nir AT plans review

YZ

  • Previous:
    • New GCS patch for John and IT#8315.
    • A long list of MODS and LBC issues/work:
      • Ported "dispMODSx2d" to Python3/Fedora.
      • Working on some MODS utilities on Fedora workstations to distinguish partner accounts.
      • Working on cleaning additional data on MODS data machines.
      • Review MODS TCS client.
      • xgterm issue on robs/x2go.
      • IT#8509 DOFPIA problem.
      • IT#8567 MODS1R scripts hang.
    • Science archive meeting (discussed MODS WCS and newdata cleaning issues).
    • TMS meeting and software work.
    • Finished safety training.
  • Next:
    • Continue to work on MODS/LBC problems.
    • TMS software test to come.
    • Other TMS work and meeting.
    • Start email thread with archive team on deploying WCS/new image utilities.

DE

  • Vault upgrade fallout minimal, most users have upgraded with no significant problems. Go users!
  • Backup changes
    • Moving to new storage
      • Local disk for regular/oops restores
      • New Synology for long term and DR
    • Need to chain backup jobs for Vault DB quiescence without breaking other jobs
    • To be clear, we are changing policy re: Vault backups from doing a native backup to an entire VM backup with Veeam DB-aware backup. Less space, less complexity, easier restores.
    • Followup again with Veeam about licensing. Kind of ridiculous process with manager approval, etc. Why are you not a web site?
  • Moved archive backup to draw from new Synology, slight mods to address path changes
  • Replacement 1TB SSD arrived
    • I think it's the RMA one rather than the ordered one. Will make sure that Cindy knows and appropriately address.
    • Office visit Wed to swap into passive unit
  • LBTI storage talk
    • Everyone feel that they are fully informed about available resources for storage both in-stock and for purchase?
    • For updated numbers as drive costs fall, we can change drive cost in the document's linked spreadsheet.
  • Tasks for this week:
    • Refinements of scripts and jobs for backups, etc. to address new Synology
    • Few more items about Vault upgrade with Ketiv, e.g. Vault password
    • Ping JW about Vault workflow
    • Ping JW about new laptop

PSG

  • Last week
    • NOMIC camera software modifications - save all FITS header items as strings.
    • Some progress, but some changes are causing crashes. Needs more work.
    • GCS changes (IT #7752), check in work in new fork?
  • This week
    • Continue working on NOMIC
    • Get GCS changes finished & checked in properly.

BM

  • Last week:
    • LBTI SW tests - everything working, made a fix to trigger skipframe button.
    • Merged testing branches (loadrecmat zero-padding fix, wrong filotto and maximum skip frame logging, drawlib free invalid drawobjects, gain ramp)
    • New HouseKeeper fix (was using the wrong opcode to send switch bcu nios read requests)
    • PB, Housekeeper, etc.: tracing BCU communication in MirrorCtrl etc.
  • This week:
    • SOUL SW tests (today)
    • ICE dispatch bug
    • ao-confcalib git (conf)
    • WFSArbitrator boot timeouts
    • Power backplace

GE

  • Last week
    • Continued testing the new RTD; see the notes in the pull requests for details.
    • Made changes to the RTD subwindowing requested by SciOps.
    • Continued work on the observer GUI; IT 6138 (problem with pauses in GUI) and IT 8529 (GUI possibly disconnecting from luci.luci).
  • This week
    • Keep testing the new RTD; see the notes in the pull requests for details.
    • Continue work on the observer GUI; IT 6138 (problem with pauses in GUI) and IT 8529 (GUI possibly disconnecting from luci.luci), and others...
    • Examine the other LUCI GUIs to check for GUI modifications from threads other than the Event Dispatch Thread.

SH

  • VMware PERC cards
  • Reconfigure AMD machine
  • AO switches delivered
  • Archive meeting
    • postfix
    • reboot archive-db1 (increase RAM and processor)
    • LBTI
  • Installing things on shell
  • Bursar's office to pay for stolen equipment
  • APC PDU
  • Moving data from old to new Synology

next:

  • Moving data from old to new Synology
  • VMware installations
    • licenses
  • Upgrade 507 computer
Topic revision: r9 - 14 Feb 2022, StephenHooper
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback